Consumption =40+0.75Y

Investment =140-10i

Government expenditure =100

Tax =80

Money deposit =0.2Y-5i

Money supply =85

(i)Find equilibrium, Y, I ?


1
Expert's answer
2021-08-24T05:54:10-0400

Equilibrium income "Y = C+1+G"

Disposable income "Yd = Y-T"

Therefore "Y= 40 + 0.75 (Y-80) + 140 - 10i + 100"

Removing the bracket we get

"Y= 40 + 0.75Y - 60 + 140 - 10i + 100"

Rearranging the equation and collecting the like terms together

We get "0.25Y = 220 - 10i"

Solving for Y by dividing both sides by 0.25

Equilibrium income is "Y* = 880 - 40i"

To work out for equilibrium interest we make i the subject of the formula

"40i= 880- Y"

dividing both sides by 40 we get

Equilibrium interest, "i* = 22 - 0.025y"